OpenStudy (anonymous):

Help me please! What is the simplest form of the number: Sqrt3(cube root 3)

OpenStudy (shamil98):

\[\sqrt[3]{3^3}\] this?

OpenStudy (shamil98):

\[or~3 \sqrt[3]{3}\]

OpenStudy (anonymous):

hmmm

OpenStudy (abmon98):

+-3

OpenStudy (anonymous):

@shamil98 The second one

OpenStudy (shamil98):

Okay. Well it's already in it's simplest form then.. unless you are asked to convert to a decimal and then multiply (the decimal is quite large).

OpenStudy (anonymous):

i thought so

OpenStudy (anonymous):

that'd b 4.something

OpenStudy (anonymous):

It's asking me to convert it to a fraction.

OpenStudy (anonymous):

0.0

OpenStudy (shamil98):

\[\huge 3*3^{\frac{1 }{ 3}}\]

OpenStudy (anonymous):

3^(1/2) 3^(1/3)=3^(5/6) Sigma Equation and LaTex is not working at this time.

OpenStudy (anonymous):

Thank you!!!

OpenStudy (shamil98):

oh you meant \[\sqrt{3} * \sqrt[3]{3}\]